OpenStudy (anonymous):

Can you show me how to solve 4y=x+1 step by step

OpenStudy (anonymous):

4y=x+1 y=(x+1)/4

OpenStudy (anonymous):

wait are you trying to solve for x? Because this isn't possible unless you know what y is. I need more info here.